S N 1 characteristics would be exhibited in the hydrolysis of a.chlorobenzene b.1-chloronapthalene c.propenyl chloride d.2-chloro 1-propene


SN1 characteristics would be exhibited in the hydrolysis of


a.chlorobenzene


b.1-chloronapthalene


c.propenyl chloride


d.2-chloro 1-propene


Grade:12th Pass

2 Answers

Aman Bansal
12 years ago

Dear Menka,

Hydrolysis of Propenyl chloride shows Sn Characteristics.

?Best Of Luck...!!!!

Cracking IIT just got more exciting,It’s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert

Abhishek Panigrahi
11 years ago

SN1 characteristics would be exhibited by propenyl chloride. 

i.e., CH2=CH-CH2-Cl + H2O → CH2=CH-CH2-OH + HCl.

This is because of greater stability of the carbocation intermediate CH2=CH-CH2+ through different resonating structures.